OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. a. y = x + 2 b. y = -x - 2 c. y = -x + 4 d. y = x - 4

OpenStudy (rob1525):

which equation when you plug in -3 for x, y = 1

OpenStudy (anonymous):

idk

OpenStudy (rob1525):

Okay, which equation has the same slope as y= -x-4

OpenStudy (anonymous):

what is a slope

OpenStudy (rob1525):

slope is rise/run

OpenStudy (anonymous):

im not under standing

OpenStudy (rob1525):

Are you just learning about this stuff?

OpenStudy (anonymous):

yes i am

OpenStudy (rob1525):

when you plug in a number for x you get a value for y right? So example : when x=5 y= -(5)-4=-9. You see?

OpenStudy (anonymous):

ok so u just sub the x with the 5

OpenStudy (rob1525):

exactlly! Now the slope is the number connected to the x and the y intercept is the other number added or subtracted from x.

OpenStudy (rob1525):

If two equations have the same slope they are said to be parallel. the y intercept is just different.

OpenStudy (anonymous):

ok so -3 = -1 - 4

OpenStudy (rob1525):

well thats -1-4 is -5

OpenStudy (anonymous):

im lost now lol

OpenStudy (rob1525):

you'er doing great just pick another equation to test.

OpenStudy (rob1525):

remember the numbers for the point is (x,y).